# Multi-Agent, Human-Agent and Beyond: A Survey on Cooperation in Social Dilemmas

### v1

@@ 3.1 Shaping Intrinsic and External Motivations
In SSDs, achieving cooperation can lead to higher collective long-term returns for all agents than each agent acting selfishly. In order to encourage agents to learn cooperative policies, it is usually necessary to introduce additional motivation rewards in addition to environmental rewards. Most works introduce the motivation by shaping agent $i$’s reward $r_{i}^{\text{total}}$ as the weighted sum of the environmental reward $r_{i}^{\text{env}}$ and the motivation reward $r_{i}^{\text{mot}}$ like the following form:
@@
where $\alpha$ and $\beta$ are weights. According to the sources of the motivation reward, we now review the mechanisms for shaping motivations in SSDs from two perspectives: the intrinsic motivation and the external motivation.
@@ 3.1.1 Intrinsic Motivation
We refer to motivations that agents not only care about one’s own but also the others’ reward as the intrinsic motivation. The idea of the intrinsic motivation in SSDs is rooted in behavioral economics and social psychology, in which research has found that individuals have the intrinsic motivation to do certain things for their inherent satisfactions rather than for some specific consequence (Ryan and Deci, 2000), e.g., inequity aversion, social value orientation, altruism, social influence, and reputation.
Inequity aversion.
Hughes et al. (2018) pioneer the incorporation of intrinsic motivation within SSDs, and define it through the lens of inequality aversion. By inequality aversion, agents prefer fairness and resist incidental inequalities. To provide a glimpse of the definition of intrinsic motivation in SSDs, we present Hughes’s formalization based on inequality aversion as follows:
@@
where $o^{i}_{t}$ is agent $i$’s observation of the global state $s_{t}$, and the agent $i$’s temporal smoothed rewards $e^{i}_{t}$ are updated by $e^{i}_{t}(o^{i}_{t},a^{i}_{t})=\gamma\lambda e^{i}_{t-1}(o^{i}_{t},a^{i}_{t})+r_{i}^{\text{env}}(o^{i}_{t},a^{i}_{t})$. Intuitively, the first term characterizes the agent $i$’s reward loss when other agents achieve rewards greater than agent $i$’s own and thus represents the disadvantage inequity aversion, and the second term characterizes the reward loss when agent $i$ performing better than others and thus represents the advantage inequity aversion. $\alpha^{i}$ and $\beta^{i}$ control agent $i$’s aversion to disadvantageous inequity and advantageous inequity. Interestingly, the authors find that advantageous inequity aversion promotes cooperation in public goods dilemmas by providing an unambiguous feedback signal, while disadvantageous inequity aversion promotes cooperation in common dilemmas by providing a noisy signal for agents’ sustainable behaviors.
Social value orientation.
Social Value Orientation (SVO) measures the magnitude of the concern people have for others. Schwarting et al. (2019) highlight the nature of SDs in autonomous driving and enhance the prediction of other drivers’ behaviors by estimating their SVOs. McKee et al. (2020) use SVO to characterize the social preferences of agents in SSDs. The relationship between the agent $i$’s own reward $r_{i}$ and the average reward of other agents $\overline{r}_{-i}$ is characterized by a reward angle $\theta(\vec{R})=\arctan\left(\overline{r}_{-i}/r_{i}\right)$. By defining the motivation reward $r_{i}^{\text{mot}}$ as $-\omega\cdot\lvert\theta^{\text{SVO}}-\theta(\vec{R})\rvert$, the agents with specific SVO given by $\theta^{\text{SVO}}$ can be trained. In addition, they also find that populations trained with heterogeneous SVO develop more universal policies with higher cooperative levels than homogeneous populations. Furthermore, Madhushani et al. (2023) find that learning best responses to diverse policies trained with heterogeneous SVO leads to better zero-shot generalization in SSDs. Li et al. (2023) propose a learning framework based on SVO called RESVO, which emerges stable roles in populations and efficiently solves SSDs through the division of labor. In addition, similar to SVO, Baker (2020) propose an environment augmentation called Randomized Uncertain Social Preferences (RUSP) to characterize social preferences, and find emergent direct reciprocity, indirect reciprocity and reputation, and team formation when training agents with RUSP.
Altruism.
@@
Altruism amounts to taking costly actions that mostly benefit other individuals. In SSDs, altruistic agents prioritize the collective reward of the entire group over their own individual rewards. Wang et al. (2019a) introduce the altruistic intrinsic motivation parameterized by a shared reward network, which is updated slowly to simulate natural selection for a higher collective reward of all agents. Their altruistic evolutionary paradigm solves difficult SSDs without handcrafting. Some works assume that individual altruism only applies to certain members within a group, which divides the entire group into different teams to promote cooperation more effectively. Radke et al. (2023) provides theoretical groundwork as to why and under which conditions smaller teams outperform larger teams.
Social influence.
@@
Many experimental studies have revealed that individual behavior is subject to social influences. Jaques et al. (2019) consider one’s social influence as the intrinsic motivation in SSDs, which is defined as to what extent the other agents change their actions because of the influencer’s action. High-level cooperation in SSDs within three social influence models demonstrates that social influence leads to enhanced coordination and communication among agents.
Reputation.
@@
Reputation serves as a crucial reciprocal mechanism for fostering cooperation within large-scale human populations. McKee et al. (2021) model one’s reputation based on the matching level of its own and group contributions in SSDs. The authors conduct MARL simulations and human behavior experiments under identifiable (contributions of each group member can be perfectly monitored) and anonymous (contributions can not be perfectly monitored) conditions. Results of several metrics of agents’ behavior demonstrate that reputation promotes cooperation among agents by developing a non-territorial, turn-taking strategy to coordinate collective action.
@@ 3.1.2 External Motivation
In contrast to intrinsic motivation, external motivation is a construct that pertains whenever an activity is done in order to attain some separable outcome (Chentanez et al., 2004). In this survey, we refer to motivation directly from other agents as external motivation.
Peer rewarding.
@@
Peer rewarding is a type of external motivation, allowing agents’ reward functions to be directly modified by others. Technically, peer rewarding improves reward sparsity and encourages agents to explore cooperative behaviors in SSDs. Lupu and Precup (2020) introduce gifting options into agents’ action space and find that zero-sum gifting solves the tragedy of the commons most effectively. Further, some work study learnable peer rewarding. Yang et al. (2020) propose the Learning to Incentivize Others (LIO) framework, where each agent learns its own incentive function by explicitly accounting for its impact on recipients’ behavior directly and its own extrinsic objective indirectly. They find that LIO agents find a near-optimal division of labor in SSDs. In addition, some economic concepts, such as market participation (Schmid et al., 2022), where agents learn to participate in others’ rewards by acquiring shares, and taxation (Hua et al., 2023), where a centralized agent learns the Pigovian tax/allowance to maximizes the social welfare, have also been introduced into SSDs to model peer rewarding.
Agreements.
@@
Human societal structures are under-girded by a multitude of agreements, which are designed to incentivize compliance through rewards for those who conform to their mandates, and to enforce discipline through penalties for those who violate the same. Can these agreements also promote cooperation among AI agents in SSDs? Vinitsky et al. (2023) construct an agent architecture called Classifier Norm Model (CNM) that can use public sanctions to spark the emergence of social norms in SSDs. The CNM agents learn to classify others’ transgression and enforce social norms from experience, and they converge on beneficial equilibria and are better at resolving free-rider problems. Christoffersen et al. (2023) consider the contracting augmentation among agents, where agents voluntarily agree to binding state-dependent transfers of reward. Gemp et al. (2022) construct D3C agents to minimize the price of anarchy, a gap between the welfare that can be achieved through perfect coordination against that achieved by self-interested agents at Nash equilibria. Eccles et al. (2019) construct online-learning reciprocal agents, who try to measure and match the level of sociality of others and influence naive agents to promote their cooperation.
@@ 3.2 Playing against Diverse Opponents
The other paradigm of solutions to SSDs is playing with diverse opponents. One may feel this resembles solving a zero-sum game at the first glance, but the difference is that in mixed-motivation SSDs, directly maximizing one’s own rewards may have opposite effects. In fact, it is more reasonable to decide one’s own choice based on the opponent’s behavior style, as is verified in the famous strategies Tit-for-Tat (TFT) or Win-Stay, Lose-Shift (WSLS) in iterated SDs. Based on the principles of constructing strategies, we divide the existing work into two categories: two-stage adaptive policy selection and online learning opponent shaping.
Adaptive policy selection.
@@
This type of method typically involves two stages: first pre-training different styles of policies; then adaptively selecting by designed strategies in real-time games. Lerer and Peysakhovich (2017) construct an approximate Markov Tit-for-Tat (amTFT) agent. They train cooperative and defective policies by reward shaping, and design a TFT-like strategy manually to decide when to cooperate or defect. Further, Wang et al. (2019b) propose the cooperation degree detection network, which is a classifier trained on trajectories generated by pre-trained policies with various degrees of cooperation. Experimental results show that the strategy can avoid being exploited by exploitative opponents and achieve cooperation with cooperative opponents. Gléau et al. (2022) introduce Circular SSD and construct Graph-based TFT for asymmetric game scenarios. Moreover, O’Callaghan and Mannion (2021) propose a method of training agents with tunable levels of cooperation in SSDs, based on multi-objective reinforcement learning.
Opponent shaping.
@@
An opponent shaping agent tries to influence the future strategies of its learning opponents online by offering feedback on their behavior. Foerster et al. (2018) propose the Learning with Opponent Learning Awareness (LOLA) rule, which takes into account the learning dynamics of the opponent. In detail, the LOLA learning rule includes an additional term that accounts for the impact of one agent’s policy on the anticipated parameter update of its opponent. The authors demonstrate that the LOLA method can achieve reciprocity-based cooperation under self-play settings by finding TFT-like strategies and also exploit other naive learning algorithms. Subsequently, a series of improvements to LOLA are proposed. Letcher et al. (2018) propose Stable Opponent Shaping that interpolates robustly between LOLA and LookAhead (Zhang and Lesser, 2010) to prevent LOLA agents from arrogant behavior and converging to non-fixed points. Willi et al. (2022) propose Consistent LOLA to address the consistency problem of LOLA. And Zhao et al. (2022) introduce Proximal LOLA which guarantees behaviorally equivalent policies result in behaviorally equivalent updates to address LOLA’s sensitivity to policy parameterization. Furthermore, in order to address the limitation of explicit gradients required by the LOLA series algorithms, Lu et al. (2022) propose Model-Free Opponent Shaping (M-FOS) that learns in a meta-game in which each meta-step is an episode of the underlying (“inner”) game. The meta-state consists of the inner policies, and the meta-policy produces a new inner policy to be used in the next episode. In addition to policy-based method, a value-based opponent shaping method called Meta-Value Learning is proposed by Cooijmans et al. (2023), which can be seen as the value-based complement to the policy gradient of M-FOS.

### v2

@@ 3.1 Shaping Intrinsic and External Motivations
In SSDs, achieving cooperation leads to higher collective long-term returns for all agents than each agent acting selfishly. In order to encourage agents to learn cooperative policies, it is usually necessary to introduce additional motivation rewards in addition to environmental rewards. Most works in this line introduce the motivation by shaping agent $i$’s reward $r_{i}^{\text{total}}$ as the weighted sum of the environmental reward $r_{i}^{\text{env}}$ and the motivation reward $r_{i}^{\text{mot}}$ like the following form:
@@
where $\alpha$ and $\beta$ are weights. According to the sources of the motivation reward, we now review the mechanisms for shaping motivations in SSDs from two perspectives: the intrinsic motivation and the external motivation. Table 1 summarizes the works cited in Section 3.1.
It is worth noting that the aim of works on mechanisms is to validate whether the introduced mechanisms can enhance cooperation in SSDs, typically by comparing the level of cooperation with and without these mechanisms. Only a few works compare the performance across different mechanisms [29, 30]. This is due to the varying assumptions and information requirements of each mechanism, which make fair comparisons challenging.
@@ 3.1.1 Intrinsic Motivation
We refer to motivations that agents not only care about one’s own but also the others’ reward as the intrinsic motivation. The idea of the intrinsic motivation in SSDs is rooted in behavioral economics and social psychology, in which research has found that individuals have the intrinsic motivation to do certain things for their inherent satisfactions rather than for some specific consequence [31], e.g., inequity aversion, social value orientation, altruism, social influence, and reputation.
Table 1: Summary of works in shaping intrinsic and external motivations in SSDs.
Motivation Mechanism Work
Intrinsic Motivation Inequity Aversion Weighted inequity aversion [Hughes et al. [8]]
Altruism Evolving shared reward network [Wang et al. [32]]
Team credo [Radke et al. [33]]
Organizational psychology[Radke et al. [34]]
Team size [Radke et al. [35]]
Social Value Orientation Driver behavior estimating with SVO[Schwarting et al. [36]]
SVO and cooperation [McKee et al. [9]]
SVO and generalization[Madhushani et al. [37]]
RESVO [Li et al. [29]]
RUSP [Baker [38]]
Social Influence Causal influence [Jaques et al. [39]]
Reputation Reputation of contribution [McKee et al. [40]]
External Motivation Peer Rewarding Gifting [Lupu and Precup [41]]
Share trading [Schmid et al. [42]]
LIO [Yang et al. [43]]
Taxation [Hua et al. [30]]
Agreement Social norm [Vinitsky et al. [44]]
Contract [Vinitsky et al. [44]]
Price of anarchy [Gemp et al. [45]]
Sociality matching [Eccles et al. [46]]
Inequity aversion
Hughes et al. [8] pioneer the incorporation of intrinsic motivation from behavioral economics and psychology explicitly within SSDs, and define it through the lens of inequality aversion. Inequity aversion is the preference for fairness and resistance to incidental inequalities [47], which consist of advantageous inequity aversion and disadvantageous inequity aversion. To provide a glimpse of the definition of intrinsic motivation in SSDs, we present Hughes’s formalization based on inequality aversion as follows:
@@
where $o^{i}_{t}$ is agent $i$’s observation of the global environmental state $s_{t}$, and the agent $i$’s temporal smoothed rewards $e^{i}_{t}$ are updated by $e^{i}_{t}(o^{i}_{t},a^{i}_{t})=\gamma\lambda e^{i}_{t-1}(o^{i}_{t},a^{i}_{t})+r_{i}^{\text{env}}(o^{i}_{t},a^{i}_{t})$. Intuitively, the first term characterizes the agent $i$’s reward loss when other agents achieve rewards greater than agent $i$’s own and thus represents the disadvantage inequity aversion, and the second term characterizes the reward loss when agent $i$ performing better than others and thus represents the advantage inequity aversion. $\alpha^{i}$ and $\beta^{i}$ control agent $i$’s aversion to disadvantageous inequity and advantageous inequity. Interestingly, the authors find that advantageous inequity aversion promotes cooperation in public goods dilemmas by providing an unambiguous feedback signal, while disadvantageous inequity aversion promotes cooperation in common dilemmas by providing a noisy signal for agents’ cooperative and sustainable behaviors.
Altruism
@@
Altruism amounts to taking costly actions that mostly benefit other individuals [48]. In SSDs, altruistic agents prioritize other agent’s reward or the collective reward of the entire group over their own individual rewards. Wang et al. [32] introduce the altruistic intrinsic motivation parameterized by a shared reward network, which is updated slowly compared to policy networks to simulate natural selection for a higher collective reward of all agents. Without handcrafting reward shaping, their altruistic evolutionary paradigm solves difficult SSDs. Some works assume that individual altruism only applies to certain members within a group, which divides the entire group into different teams to promote cooperation more effectively. [33, 34]. Radke et al. [35] discuss advantages and disadvantages of team division on full-group cooperation. They find the optimal size for team division to promote cooperation in the entire group and prove it both theoretically and experimentally.
Social value orientation
Social Value Orientation (SVO) measures how an agent apportion rewards of itself and others in its reward function[49, 50]. Aforementioned inequity aversion and altruism can be consider as particular cases of SVO. Here we focus on works that study impact of general SVO rather than considering only specific orientation. Schwarting et al. [36] highlight the nature of SDs in autonomous driving and enhance the prediction of other drivers’ behaviors by estimating their SVOs. McKee et al. [9] first use SVO to characterize the social preferences of agents in SSDs. The relationship between the agent $i$’s own reward $r_{i}$ and the average reward of other agents $\overline{r}_{-i}$ is characterized by a reward angle $\theta(\vec{R})=\arctan\left(\overline{r}_{-i}/r_{i}\right)$. By defining the motivation reward $r_{i}^{\text{mot}}$ as $-\omega\cdot\lvert\theta^{\text{SVO}}-\theta(\vec{R})\rvert$, the agents with specific SVO given by $\theta^{\text{SVO}}$ can be trained. In addition, they also find that populations trained with heterogeneous SVO develop more universal policies with higher levels of cooperation than homogeneous populations. Furthermore, Madhushani et al. [37] find that learning best responses to diverse policies trained with heterogeneous SVO leads to better zero-shot generalization in SSDs. Li et al. [29] propose a learning framework based on SVO called RESVO, which emerges stable roles in populations and efficiently solves SSDs through the division of labor. In addition, similar to SVO, Baker [38] propose an environment augmentation called Randomized Uncertain Social Preferences (RUSP) to characterize social preferences, and find emergent direct reciprocity, indirect reciprocity and reputation, and team formation when training agents with RUSP.
Social influence
@@
Many experimental studies have revealed that individual behavior is subject to social influences [51, 52]. Jaques et al. [39] introduce social influencers into SSDs, who use social influence as their intrinsic reward. The social influence is defined as a causal influence: to what extent the other agents change their actions because of the influencer’s action. This is similar to the concept of informational social influence, which is defined as an influence to accept information obtained from another as evidence about the nature of the game [53]. High level of cooperation in SSDs within three social influence models encompassing basic influence model, influential communication and modeling other agents demonstrates that introducing of social influencers leads to enhanced coordination and communication among agents.
Reputation
@@
Reputation serves as a crucial reciprocal mechanism for fostering cooperation within large-scale human populations [54]. McKee et al. [40] model one’s reputation based on the matching level of its own and group contributions in SSDs. The authors conduct MARL simulations and human behavior experiments under identifiable (contributions of each group member can be perfectly monitored) and anonymous (contributions can not be perfectly monitored) conditions. Results of several metrics of agents’ behavior demonstrate that reputation promotes cooperation among agents by developing a non-territorial, turn-taking strategy to coordinate collective action.
@@ 3.1.2 External Motivation
In contrast to intrinsic motivation, external motivation is a construct that pertains whenever an activity is done in order to attain some separable outcome [55]. In this survey, we refer to motivation directly from other agents as external motivation.
Peer rewarding
@@
Peer rewarding is a type of external motivation, allowing agents’ reward functions to be directly modified by others. Technically, peer rewarding improves reward sparsity and encourages agents to explore cooperative behaviors in SSDs. Lupu and Precup [41] introduce gifting options into agents’ action space and find that zero-sum gifting solves the tragedy of the commons most effectively. They introduce three types of gifting: zero-sum, where agents can send a gift $g$ but incur an immediate penalty of $-g$ from its gifting; fixed budget, where agent can gift $g$ from a fixed budget $B$ in each episode; and replenishable budget, where gift budget can increment as the quantity of rewards it can collect from the environment. Further, some work study learnable peer rewarding. Yang et al. [43] propose the Learning to Incentivize Others (LIO) framework, where each agent learns its own incentive function by explicitly accounting for its impact on recipients’ behavior directly and its own extrinsic objective indirectly. They find that LIO agents find a near-optimal division of labor in SSDs. In addition, some economic concepts, such as market participation [42], where agents learn to participate in others’ rewards by acquiring shares, and taxation [30], where a centralized agent learns the Pigovian tax/allowance to maximizes the social welfare, have also been introduced into SSDs to model peer rewarding.
Agreement
@@
Human societal structures are under-girded by a multitude of agreements, which are designed to incentivize compliance through rewards for those who conform to their mandates, and to enforce discipline through penalties for those who violate the same. Can these agreements also promote cooperation among AI agents in SSDs? Vinitsky et al. [44] construct an agent architecture called Classifier Norm Model (CNM) that can use public sanctions to spark the emergence of social norms in SSDs. The CNM agents learn to classify others’ transgression and enforce social norms from experience, and they converge on beneficial equilibrium and are better at resolving free-rider problems. Christoffersen et al. [11] consider the contracting augmentation among agents, where agents voluntarily agree to binding state-dependent transfers of reward. Gemp et al. [45] construct D3C agents to minimize the price of anarchy, a gap between the welfare that can be achieved through perfect coordination against that achieved by self-interested agents at Nash equilibrium. D3C resembles the celebrated Win-Stay-Lose-Shift [56] strategy to improve inefficient equilibrium in SSDs. Eccles et al. [46] construct online-learning reciprocal agents, who try to measure and match the level of sociality of others and influence naive agents to promote their cooperation. The results indicate that reciprocal agents can influence naive agents to promote their cooperation.
Table 2: Summary of works in playing against diverse opponents in SSDs.
Category Work
Adaptive Policy Adopting amTFT [Lerer and Peysakhovich [12]]
Cooperation degree detection [Wang et al. [57]]
Graph-based TFT [Gléau et al. [58]]
Tunable agent [O’Callaghan and Mannion [59]]
Opponent Shaping LOLA [Foerster et al. [13]]
SOS [Letcher et al. [60]]
COLA [Willi et al. [61]]
POLA [Zhao et al. [62]]
M-FOS [Lu et al. [63]]
Meta-Value Learning [Cooijmans et al. [64]]
@@ 3.2 Playing against Diverse Opponents
The other paradigm of solutions to SSDs is playing with diverse opponents. One may feel this resembles solving a zero-sum game at the first glance, but the difference is that in mixed-motivation SSDs, directly maximizing one’s own rewards may have opposite effects. In fact, it is more reasonable to decide one’s own choice based on the opponent’s behavior style, as is verified in the famous strategies Tit-for-Tat (TFT) or Win-Stay-Lose-Shift (WSLS) in iterated SDs. Based on the principles of constructing strategies, we divide the prevailing work into two categories: two-stage adaptive policy adopting and online learning opponent shaping. Table 2 summarizes the works cited in Section 3.2.
Adaptive policy adopting
@@
This type of method typically involves two stages: first pre-training different styles of policies; then adaptively selecting by designed strategies in real-time games. Lerer and Peysakhovich [12] construct an approximate Markov Tit-for-Tat (amTFT) agent. They train cooperative and defective policies by reward shaping, and design a TFT-like strategy manually to decide when to cooperate or defect. Further, Wang et al. [57] propose the cooperation degree detection network, which is a classifier trained on trajectories generated by pre-trained policies with various degrees of cooperation. Experimental results show that the strategy can avoid being exploited by exploitative opponents and achieve cooperation with cooperative opponents. Gléau et al. [58] introduce Circular SSD and construct Graph-based TFT for asymmetric game scenarios. Moreover, O’Callaghan and Mannion [59] propose a method of training agents with tunable levels of cooperation in SSDs, based on multi-objective reinforcement learning.
Opponent shaping
An opponent shaping agent tries to influence the future strategies of its learning opponents online by offering feedback on their behavior. Foerster et al. [13] propose the Learning with Opponent Learning Awareness (LOLA) rule, which takes into account the learning dynamics of the opponent. In detail, the LOLA learning rule includes an additional term that accounts for the impact of one agent’s policy on the anticipated parameter update of its opponent:
$V^{1}(\theta^{1},\theta^{2}+\Delta\theta^{2})\approx V^{1}(\theta^{1},\theta^{2})+(\Delta\theta^{2})^{T}\nabla_{\theta^{2}}V^{1}(\theta^{1},\theta^{2}),$
@@
where $\nabla$ is the gradient operator, $V^{i}(s;\theta^{1},\theta^{2})$ is the state-value function (expected cumulative rewards with the initial state $s$) of agent $i$ parameterized by two agents’ policy parameters $\theta^{1}$ and $\theta^{2}$, and the update of $\theta^{2}$ follows $\Delta\theta^{2}=\eta\nabla_{\theta^{2}}V^{2}(s;\theta^{1},\theta^{2})$ with $\eta>0$ being the step size. The authors demonstrate that LOLA can achieve reciprocity-based cooperation under self-play settings by finding TFT-like strategies and also exploit other naive learning algorithms. Subsequently, a series of improvements to LOLA are proposed. Letcher et al. [60] propose Stable Opponent Shaping that interpolates robustly between LOLA and LookAhead [65] to prevent LOLA agents from arrogant behavior and converging to non-fixed points. Willi et al. [61] propose Consistent LOLA to address the consistency problem of LOLA. And Zhao et al. [62] introduce Proximal LOLA which guarantees behaviorally equivalent policies result in behaviorally equivalent updates to address LOLA’s sensitivity to policy parameterization. Furthermore, in order to address the limitation of explicit gradients required by the LOLA series algorithms, Lu et al. [63] propose Model-Free Opponent Shaping (M-FOS) that learns in a meta-game in which each meta-step is an episode of the underlying (“inner”) game. The meta-state consists of the inner policies, and the meta-policy produces a new inner policy to be used in the next episode. In addition to policy-based method, a value-based opponent shaping method called Meta-Value Learning is proposed by Cooijmans et al. [64], which can be seen as the value-based complement to the policy gradient of M-FOS.
Table 3: Summary of works in human-agent cooperation and leveraging agents to motivate human-human cooperation in SDs.
Subject Motivation Work
Human-agent Algorithms design S++ and S# [Crandall [66], Crandall et al. [14], Oudah et al. [67]]
cooperation Exploitation avoiding Zero-determinant strategy [Hilbe et al. [15], Wang et al. [68]]
Safety control [Zhang et al. [69]]
Experiment-driven insights Partner selection [Santos et al. [70, 71]]
Human bias towards agents Transparency-efficiency tradeoff [Ishowo-Oloko et al. [16]]
Algorithm exploitation [Karpus et al. [72]]
Guilt and envy [Melo et al. [73]]
Human bias mitigating Culture and emotion [de Melo and Terada [17]]
Verbal communication [Maggioni and Rossignoli [74]]
Human-human Network engineering Social planner [Shirado and Christakis [75], McKee et al. [18]]
cooperation Evolutinary dynamics Fixed behavior [Terrucha et al. [76], Guo et al. [19], Sharma et al. [77]]
Opportunity cost [Han et al. [78]]
Authority grant Considering long-term profit [de Melo et al. [79]]
Predefined and customized agents [Fernández Domingos et al. [80]]